For a sample of size 36, ∑x = 761.6, ∑x2 = 16125.5. Is the population mean 21


1
Expert's answer
2021-04-15T06:49:02-0400

mean = "\frac{\sum{x}}{sample \ size}" = "\frac{761.6}{36} \approx 21"

Answer: yes, mean = 21.